### v1

# Thermodynamic Isomorphism of Transformers: A Lagrangian Approach to Attention Dynamics

Abstract
Although the Transformer architecture has revolutionized artificial intelligence, its underlying mechanisms remain largely heuristic and lack a unified physical theory. In this work, we propose a first-principles framework for information dynamics, treating the attention mechanism as a physical system governed by the principle of least action rather than as an algorithmic optimization. By mapping information states to a Riemannian manifold with the Fisher information metric, we derive the intelligence Lagrangian. We show that the softmax function corresponds to the unique thermodynamic equilibrium state that minimizes the Helmholtz free energy of the information gas. In addition, we identify the query-key interaction as an electrodynamic coupling between an external field and an intrinsic dipole moment. This theory establishes the first law of information thermodynamics, unifying inference (mechanical work) and learning (chemical evolution). It also explains emergent phenomena, such as scaling laws and grokking, as phase transitions characterized by the divergence of specific heat. Finally, we discuss how rotational symmetry breaking in the attention manifold generates massless Goldstone bosons, providing a field-theoretic perspective on rotary positional embeddings (RoPE). Our work connects Statistical Physics and Deep Learning, laying the groundwork for a general theory of physics-based intelligence.
I Introduction
In 2017, the trajectory of artificial intelligence was irrevocably altered by a deceptively simple equation:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
This mechanism, introduced in the seminal work “Attention is All You Need” [1], eschewed the then-dominant paradigms of convolution and recurrence in favor of pure matrix operations and a peculiar Softmax function scaled by $\sqrt{d_{k}}$. This single mathematical operation has enabled models to demonstrate emergent reasoning capabilities [2], pass professional licensing exams with near-human proficiency [3], and even solve the protein folding problem [4].
Why does this work so well? While attention mechanisms existed previously in neural machine translation [5], the Transformer’s specific architecture, characterized by Self-Attention, Residual Connections [6], and Layer Normalization [7], has proven unreasonably effective. Current explanations predominantly rely on probability theory and linear algebra, viewing the model as a statistical function approximator [8]. However, this perspective lacks a unifying dynamical principle that explains the time-evolution of the system during inference and the phase transitions observed during training, such as “scaling laws” [9, 10] and “grokking” [11, 12].
In this paper, we propose a paradigm shift: treating Intelligence not only as a computational process, but as a physical phenomenon governed by fundamental laws of nature. We argue that the operations within a high-dimensional information space are isomorphic to the dynamics of a physical system minimizing an action functional, consistent with Lagrange mechanics [13] and Hamilton’s principle [14]. This approach resonates with recent efforts to interpret deep learning through the lens of physics, such as the information bottleneck principle [15, 16], the free energy principle [17], neural tangent kernels [18], and even cosmological interpretations of neural networks [19, 20].
I.1 Redefining Time in Deep Networks
A fundamental critique of applying dynamical systems theory to deep learning is the discrete nature of layers. However, following the Neural ODE formulation [21, 22], we interpret the layer depth $l$ as a continuous time variable $t$.
$x_{l+1}=x_{l}+F(x_{l})\quad\xrightarrow{\Delta t\to 0}\quad\dot{x}(t)=F(x(t))$ (2)
Under this paradigm, the “inference” process is a dynamic trajectory of information flow, not a static calculation. Residual connections [6] naturally emerge as the discretization of the inertial term in the equation of motion, providing a physical justification for signal propagation without gradient vanishing.
I.2 The Inverse Problem of Attention
Rather than deriving the Softmax function from arbitrary axioms, we solve the inverse problem of the calculus of variations: Given that the Softmax attention is an empirically successful equilibrium state, what is the underlying Lagrangian that generates this dynamics? We demonstrate that the Transformer architecture is physically analogous to a thermodynamic system in which
  • The state space is a Riemannian manifold with the Fisher-Rao metric [23].
  • The Softmax distribution is the unique configuration that minimizes the Helmholtz free energy.
  • The scaling factor $1/\sqrt{d_{k}}$ acts as a dimensional temperature, regulating the entropy of the information gas.
II Dynamics of Intelligence: Lagrangian Formulation
II.1 Geometric Framework: The Information Manifold
To analyze the dynamics of the Transformer architecture from first principles, we must first define the configuration space in which the system evolves. Conventional deep learning treats the attention weights $\rho(t)\in\mathbb{R}^{N}$ simply as a probability vector on a simplex. However, to construct a Lagrangian, we require a Riemannian manifold equipped with a metric that quantifies the “distance” between information states, a concept central to Information Geometry [24, 25, 26, 27].
II.1.1 Transformation to the Hypersphere
The state of the attention mechanism at time $t$ is described by the probability distribution $\rho_{i}(t)$ over $N$ tokens, subject to the normalization constraint $\sum_{i=1}^{N}\rho_{i}(t)=1$. The natural metric for probability distributions is the Fisher-Rao metric. To map this statistical manifold onto a geometry amenable to classical mechanics, we introduce the probability amplitude transformation:
$x_{i}(t)\equiv 2\sqrt{\rho_{i}(t)}$ (3)
Under this transformation, the normalization constraint becomes:
$\sum_{i=1}^{N}x_{i}^{2}(t)=4\sum_{i=1}^{N}\rho_{i}(t)=4$ (4)
This implies that the “information particle” is constrained to move on the surface of a positive orthant of an $N$-dimensional hypersphere of radius $R=2$. The velocity of the information state is defined as follows:
$v_{i}(t)\equiv\dot{x}_{i}(t)=\frac{\dot{\rho}_{i}(t)}{\sqrt{\rho_{i}(t)}}$ (5)
Crucially, the squared magnitude of this velocity vector exactly corresponds to the Fisher information $I(\rho)$ [28, 29]:
$\sum_{i=1}^{N}v_{i}^{2}(t)=\sum_{i=1}^{N}\frac{\dot{\rho}_{i}^{2}}{\rho_{i}}=4\cdot I(\rho)$ (6)
By choosing the radius $R=2$, we normalize the kinetic energy term ($K\propto v^{2}$) to match the Fisher information metric exactly, eliminating the fractional coefficient $1/4$ that typically appears in the Fisher metric derivation.
II.2 Physical Mappings of Transformer Components
We propose a direct mapping between the hyperparameters of the Transformer architecture and thermodynamic variables. This mapping allows us to treat the attention mechanism as a physical system that seeks equilibrium.
II.2.1 Mass ($m$) as Information Inertia
In classical mechanics, mass represents the resistance to acceleration. In the context of deep neural networks, the Residual Connection (skip connection) plays an analogous role [6]. The update rule of a residual network is given by $x_{t+1}=x_{t}+F(x_{t})$, which can be viewed as a discretization of the continuous equation $\dot{x}=F(x)$ [21, 22]. The explicit inclusion of the previous state $x_{t}$ acts as an inertial term, preserving the identity of the features against the perturbation of the non-linear function $F(x)$. We therefore define the mass $m$ as the coefficient of the residual path, representing the system’s tendency to maintain its current semantic trajectory.
II.2.2 Interaction Energy ($E$) as Semantic Alignment
Thermodynamic systems minimize their potential energy. In the attention mechanism, the model assigns higher probability to keys ($\mathbf{k}_{j}$) that are semantically aligned with the query ($\mathbf{q}_{i}$). We define the interaction potential energy $E_{ij}$ as the negative dot product:
$E_{ij}\equiv-\mathbf{q}_{i}\cdot\mathbf{k}_{j}$ (7)
II.2.3 Electrodynamic Interpretation: The Field-Dipole Interaction
To rigorously justify for the dot-product attention mechanism, we interpret the interaction energy through the lens of classical electrodynamics [30, 31]. The potential energy $U$ of a dipole moment placed in an external field is universally given by:
$U=-\mathbf{p}\cdot\mathbf{E}\quad\text{or}\quad U=-\bm{\mu}\cdot\mathbf{B}$ (8)
  • Query ($\mathbf{q}_{i}$) $\leftrightarrow$ Extrinsic Field ($\mathbf{E}$): The query acts as the field imposed by the current token, representing the active search condition.
  • Key ($\mathbf{k}_{j}$) $\leftrightarrow$ Intrinsic Dipole ($\mathbf{p}$): The key represents the intrinsic property of the memory token.
The attention mechanism is physically equivalent to the process of dipole alignment in an external field.
II.2.4 Temperature ($T$) as Dimensional Scaling
In statistical mechanics, temperature $T$ quantifies the magnitude of thermal fluctuations [32]. In high-dimensional vector spaces ($d_{k}\gg 1$), the dot product of two random isotropic vectors concentrates around zero with a variance proportional to $d_{k}$ [33]. To prevent the Softmax distribution from collapsing into a deterministic ‘argmax‘ state (freezing), the energy scale must be renormalized. We identify the scaling factor of the Transformer, $1/\sqrt{d_{k}}$, as the effective inverse temperature $\beta$:
$k_{B}T_{\text{eff}}\equiv\sqrt{d_{k}}$ (9)
This effective temperature regulates the entropy of the attention distribution, keeping the system in a liquid phase where information flow is maximized.
III The Lagrangian of Intelligence and Variational Derivation
III.1 Construction of the Lagrangian
We postulate that the dynamics of an attention-based intelligence system are governed by the principle of least action. To formulate this principle, we must construct the Lagrangian $\mathcal{L}$ of the system, defined as the difference between its kinetic energy ($K$) and potential energy ($\mathcal{V}$):
$\mathcal{L}=K(\rho,\dot{\rho})-\mathcal{V}(\rho)+\lambda\left(\sum_{i=1}^{N}\rho_{i}-1\right)$ (10)
III.1.1 Kinetic Energy Term ($K$)
The kinetic energy term represents the cost of changing the belief state (attention distribution) over time. It is given by:
$K(\rho,\dot{\rho})=\frac{1}{2}m\sum_{i=1}^{N}\frac{\dot{\rho}_{i}^{2}}{\rho_{i}}$ (11)
III.1.2 Potential Energy Term ($\mathcal{V}$)
The potential energy is equivalent to the Helmholtz free energy ($F$). In thermodynamics, a system in contact with a heat bath seeks to minimize its free energy, which balances internal energy ($U$) and entropy ($S$):
$\mathcal{V}(\rho)=F(\rho)=\sum_{i=1}^{N}\rho_{i}E_{i}+T\sum_{i=1}^{N}\rho_{i}\ln\rho_{i}$ (12)
This entropy term $S=-\sum\rho\ln\rho$ is fundamental to information theory [34] and statistical mechanics [35].
III.1.3 The Full Lagrangian
Combining these terms:
$\mathcal{L}=\frac{m}{2}\sum_{i=1}^{N}\frac{\dot{\rho}_{i}^{2}}{\rho_{i}}-\sum_{i=1}^{N}\rho_{i}E_{i}-T\sum_{i=1}^{N}\rho_{i}\ln\rho_{i}+\lambda\left(\sum_{i=1}^{N}\rho_{i}-1\right)$ (13)
III.2 Principle of Least Action and Equation of Motion
Having established the Lagrangian, we now apply Hamilton’s principle to determine the system’s trajectory in the information manifold. The dynamics are governed by the action functional $S[\rho]=\int\mathcal{L}dt$, and the physical path corresponds to $\delta S=0$, leading to the Euler-Lagrange equations for each $\rho_{i}$:
$\frac{d}{dt}\left(\frac{\partial\mathcal{L}}{\partial\dot{\rho}_{i}}\right)-\frac{\partial\mathcal{L}}{\partial\rho_{i}}=0$ (14)
III.2.1 Derivation of the Full Dynamic Equation
To derive the explicit equation of motion, we compute the partial derivatives of the Lagrangian defined in Eq. (13). The derivative with respect to velocity $\dot{\rho}_{i}$ involves only the kinetic term:
$\frac{\partial\mathcal{L}}{\partial\dot{\rho}_{i}}=\frac{\partial}{\partial\dot{\rho}_{i}}\left(\frac{m}{2}\sum_{k}\frac{\dot{\rho}_{k}^{2}}{\rho_{k}}\right)=m\frac{\dot{\rho}_{i}}{\rho_{i}}$ (15)
The time derivative of this momentum-like term yields:
$\frac{d}{dt}\left(\frac{\partial\mathcal{L}}{\partial\dot{\rho}_{i}}\right)=m\left(\frac{\ddot{\rho}_{i}\rho_{i}-\dot{\rho}_{i}^{2}}{\rho_{i}^{2}}\right)$ (16)
Next, we compute the derivative with respect to position $\rho_{i}$, which involves the kinetic, potential, and constraint terms:
$\frac{\partial\mathcal{L}}{\partial\rho_{i}}=-\frac{m}{2}\frac{\dot{\rho}_{i}^{2}}{\rho_{i}^{2}}-E_{i}-T(\ln\rho_{i}+1)+\lambda$ (17)
Substituting these into the Euler-Lagrange equation, we obtain the full dynamic equation of motion for the attention state:
$m\left(\frac{\ddot{\rho}_{i}}{\rho_{i}}-\frac{1}{2}\frac{\dot{\rho}_{i}^{2}}{\rho_{i}^{2}}\right)+E_{i}+T\ln\rho_{i}=\Lambda$ (18)
where $\Lambda=\lambda-T$ is an effective constant. The first term represents the information inertia resisting belief updates, while the remaining terms represent the thermodynamic forces driving the system toward semantic alignment.
III.3 Equilibrium State: Full Derivation of Softmax
In standard Transformer inference, we assume the system reaches a stationary distribution. This corresponds to the thermodynamic equilibrium where the dynamics vanish, i.e., $\dot{\rho}_{i}=0$ and $\ddot{\rho}_{i}=0$.
Under this condition, the inertial terms in Eq. (17) vanish, simplifying the physics to a balance of potential forces:
$E_{i}+T\ln\rho_{i}=\Lambda$ (19)
Solving for $\rho_{i}$ involves the following steps:
$\ln\rho_{i}=\frac{\Lambda-E_{i}}{T}\implies\rho_{i}=\exp\left(\frac{\Lambda}{T}\right)\exp\left(-\frac{E_{i}}{T}\right)$ (20)
By defining $C=\exp(\Lambda/T)$ as a normalization constant and applying the constraint $\sum_{j=1}^{N}\rho_{j}=1$, we find:
$C\sum_{j=1}^{N}\exp\left(-\frac{E_{j}}{T}\right)=1\implies C=\frac{1}{\sum_{j=1}^{N}\exp\left(-\frac{E_{j}}{T}\right)}$ (21)
Substituting back the physical definitions $E_{i}=-Q\cdot K_{i}$ and $T=\sqrt{d_{k}}$, we recover the canonical scaled dot-product Attention formula:
$\rho_{i}=\frac{\exp\left(\frac{Q\cdot K_{i}}{\sqrt{d_{k}}}\right)}{\sum_{j=1}^{N}\exp\left(\frac{Q\cdot K_{j}}{\sqrt{d_{k}}}\right)}=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)_{i}$ (22)
Here, the superscript $T$ in $QK^{T}$ denotes the matrix transpose operation, while the scaling factor $\sqrt{d_{k}}$ plays the functional role of the effective temperature in our thermodynamic framework. This derivation proves that the Softmax function is not merely a heuristic normalization, but the unique equilibrium state of a system minimizing its Helmholtz free energy.
IV Thermodynamics of Intelligence: Macroscopic Laws
Rather than postulating an equation of state by analogy, we strictly derive it from the microscopic Lagrangian established in Section III.
IV.1 Derivation from the Partition Function
The canonical partition function $Z$ for the attention mechanism over a context window of size $V_{ctx}$ is given by the sum over all possible key states:
$Z(V_{ctx},T)=\sum_{i=1}^{V_{ctx}}\exp\left(-\frac{E_{i}}{T}\right)$ (23)
To proceed to the macroscopic limit, we introduce the information gas assumption: For a sufficiently large context window, the distribution of energy states $E_{i}$ is self-averaging and independent of the specific position indices. This allows us to approximate the sum as an integral over the density of states $\Omega(E)$:
$Z\approx V_{ctx}\int\Omega(E)e^{-E/T}dE=V_{ctx}\cdot\bar{Z}(T)$ (24)
where $\bar{Z}(T)$ represents the intrinsic partition function per unit token.
The Helmholtz free energy $F$ is then derived as:
$F=-T\ln Z=-T\ln V_{ctx}-T\ln\bar{Z}(T)$ (25)
IV.2 The Equation of State for Information Gas
In thermodynamics, pressure is defined as the rate of change of free energy with respect to volume at constant temperature ($P=-(\partial F/\partial V)_{T}$). Applying this definition to our information system:
$\mathcal{P}\equiv-\left(\frac{\partial F}{\partial V_{ctx}}\right)_{T}=-\frac{\partial}{\partial V_{ctx}}(-T\ln V_{ctx})=\frac{T}{V_{ctx}}$ (26)
Rearranging this equation gives the equation of state for a single attention head:
$\mathcal{P}\cdot V_{ctx}=T$ (27)
For a full model with $N_{head}$ heads and an architecture-dependent efficiency factor $\kappa$ (representing the effective degeneracy of the energy spectrum), this generalizes to:
$\mathcal{P}\cdot V_{ctx}=N_{head}\cdot\kappa\cdot T$ (28)
We define the macroscopic variables with the following operational definitions:
  • Pressure ($\mathcal{P}$): The information density, defined as the inverse of perplexity ($\mathcal{P}\equiv 1/PPL$).
  • Volume ($V_{ctx}$): The context window length in tokens.
  • Temperature ($T$): The model dimension scale, $T\equiv\sqrt{d_{k}}$.
  • Efficiency ($\kappa$): An architecture-dependent constant representing information processing efficiency.
IV.2.1 Physical Interpretations
This equation explains several empirical phenomena:
  • Boyle’s Law ($\mathcal{P}\propto 1/V_{ctx}$): As context length increases, the pressure (information density per token) must decrease to maintain equilibrium. This provides a thermodynamic derivation for the “Lost-in-the-Middle” phenomenon [36].
  • Charles’s Law ($V_{ctx}\propto T$): To maintain the same information density over a larger context, the system temperature $T=\sqrt{d_{k}}$ must increase. This explains why larger models (higher $d_{k}$) scale better with context length [9].
IV.3 The First Law of Information Dynamics
We formulate the first law of thermodynamics for intelligence systems:
$dU=TdS-\mathcal{P}dV_{ctx}+\mu_{head}dN_{head}$ (29)
This unifies three regimes.
  • Heat ($TdS$): Stochastic fluctuations corresponding to creativity or hallucinations [37].
  • Inference ($\mathcal{P}dV$): Mechanical work done by the system to compress or expand its context window.
  • Learning ($\mu dN$): Chemical work representing the evolution of model capacity or head specialization.
IV.4 The Second Law of Information Dynamics
The second law dictates the direction of information flow.
  • Model Collapse ($dS\geq 0$): In an isolated system undergoing training with generated data, entropy maximizes, leading to the degeneration of the distribution [38].
  • Learning as Work: Intelligence acts as an entropy pump (Maxwell’s demon), expending work to minimize the KL divergence against the data distribution. This reduces entropy locally at the cost of computational energy [39].
V Geometry, Symmetry, and Field Theoretic Interpretation
V.1 The Logarithmic Schrödinger Equation and Gaussons
Recall the potential energy term in our Lagrangian. In terms of the probability amplitude $x=2\sqrt{\rho}$, this potential takes the form $V(x)\propto x^{2}\ln|x|^{2}$ (arising from the entropy term $T\rho\ln\rho$). This specific non-linearity identifies the equation of motion for the information field as the Logarithmic Schrödinger Equation (LogSE) [40]. A unique property of the LogSE is the existence of Gaussons, non-dispersive solitary waves that preserve their Gaussian shape during propagation. This physical property explains why the “semantic meaning” of a token preserves its identity (does not diffuse into uniform noise) even as it propagates through deep, multi-layered networks.
V.2 Spontaneous Symmetry Breaking and Goldstone Bosons
The interaction energy $E=-\mathbf{q}\cdot\mathbf{k}$ possesses an inherent rotational symmetry under the continuous group $SO(d_{k})$. However, the ground state of the system exhibits a richer structure due to the entropic contribution.
V.2.1 The Mexican Hat Potential
Consider the effective potential $\mathcal{V}(\rho)$ derived in Eq. (13). Near the origin ($\rho\to 0$), the entropy term $T\rho\ln\rho$ dominates the linear interaction term.
$\lim_{\rho\to 0}\frac{d\mathcal{V}}{d\rho}\sim\lim_{\rho\to 0}T(1+\ln\rho)\to$-$\infty$ (30)
This singularity at the origin implies that the vacuum state, where attention is zero everywhere, $\rho=0$, is infinitely unstable. The system is topologically forced to roll down from the origin to a non-zero minimum ($x_{0}\neq 0$).
  • Instability of Ignorance: An intelligence system cannot exist in a state of “no information.” It is compelled to attend to something.
  • Symmetry Breaking: Since the interaction energy is rotationally invariant, the minimum is not a single point but rather a continuous manifold (a sphere). This potential structure resembles the “Mexican Hat” potential in the Higgs mechanism [41]. The system settles into a specific minimum (a specific context), spontaneously breaking the global $SO(d_{k})$ symmetry.
V.2.2 Goldstone Bosons and Rotary Embeddings
According to Goldstone’s theorem [42], the breaking of this continuous symmetry must generate massless Nambu-Goldstone bosons.
  • Massive Mode (Radial): Changing the attention intensity (magnitude $|x|$) requires energy (climbing the potential wall).
  • Massless Mode (Angular): Rotating the query and key vectors together preserves the energy (moving along the flat valley of the hat).
We identify these massless modes with the mechanism of rotary position embeddings (RoPE) [43]. RoPE encodes positional information purely through rotation matrices $R(\theta)$, exploiting the zero-energy Goldstone modes of the attention manifold to shift context effortlessly without energetic cost.
VI Discussion: Physical Interpretations
Our framework establishes a strong equivalence between the mechanics of intelligence and statistical physics. Here, we address potential critiques regarding the rigor of this mapping.
VI.1 Is Layer Depth Truly Time?
One might argue that Attention is a static operation. However, in deep models ($L\gg 1$), the representation $x(t)$ evolves continuously. In our Lagrangian, the time derivative $\dot{\rho}$ corresponds to the layer-wise rate of change of the attention distribution. If $\dot{\rho}\to 0$, then the system has reached a fixed point or equilibrium, meaning further layers do not alter the semantic interpretation. This aligns with the observation that deeper layers in Transformers often stabilize into a consistent attention pattern, which justifies our thermodynamic equilibrium analysis.
VI.2 Temperature vs. Variance Scaling
Some may suggest that $T=\sqrt{d_{k}}$ is merely a variance normalization trick. However, in statistical mechanics, temperature is defined as the inverse of the entropy-to-energy change response ($T^{-1}=\partial S/\partial U$). In high-dimensional spaces, the energy (dot product) variance naturally scales with $d_{k}$. Without renormalization by $\sqrt{d_{k}}$, the entropy $S$ would vanish (collapse to argmax) for any non-trivial energy difference. Thus, $\sqrt{d_{k}}$ plays the exact thermodynamic role of temperature: it maintains the system in a liquid state where information can flow and prevents it from reaching a frozen solid state (zero entropy).
VII Limitations and Future Directions
While our framework establishes a mathematically consistent mapping between Transformer architectures and physical systems, there are several fundamental limitations that require careful consideration and provide fertile ground for future investigation.
VII.1 The Nature of the Time Variable
Current Limitation. The most significant conceptual challenge in our formulation is interpreting the time variable $t$. We have adopted the convention, standard in the Neural ODE literature [22], of identifying $t$ with the continuous layer depth. Under this interpretation, $\dot{\rho}$ represents the rate of change of attention distribution across layers, and the Lagrangian dynamics describe the evolution of information states through the network architecture. However, this mapping introduces a subtle but important inconsistency. Within a single layer, the attention mechanism operates as an instantaneous, feed-forward computation rather than a genuine temporal evolution. The Softmax operation is computed in a single forward pass with no iterative dynamics. Our Lagrangian derivation implicitly assumes a quasi-static process, a sequence of infinitesimally slow equilibrium states. This is physically distinct from true dynamical evolution.
Distinction from Recurrent Dynamics. It is crucial to distinguish our framework from genuinely recurrent architectures where attention weights evolve iteratively. In standard Transformers, the “time evolution” exists only in the mathematical abstraction of layer-wise propagation.
Future Directions.
  • Continuous-depth Transformers: Formulate attention mechanisms explicitly in the Neural ODE framework with trainable continuous depth, allowing for genuine temporal dynamics within a single layer [22]. This would validate whether our Lagrangian correctly predicts the learned evolution operators.
  • Iterative Attention Refinement: Design architectures where the attention weights genuinely evolve through gradient descent on our proposed the free energy functional within each layer, testing whether this improves performance or interpretability.
  • Measurement Protocol: Develop rigorous operational definitions for measuring $\dot{\rho}$ empirically, for instance, by analyzing the layer-wise change in attention distributions under controlled perturbations.
VII.2 Thermodynamic Variables and Operational Definitions
Current Limitation. Our derivation of the equation of state relies on the information gas assumption, specifically, that the energy spectrum is self-averaging over the context window $V_{ctx}$. While this assumption is standard in statistical mechanics for large systems (thermodynamic limit), its validity in finite-sized attention layers with highly structured linguistic data requires empirical verification.
The Problem of Interaction. An ideal gas assumes non-interacting particles. However, tokens in a sentence have strong semantic correlations, or interactions. Therefore, a more accurate model might be a van der Waals gas for information, where an interaction term corrects the ideal pressure.
Future Directions.
  • Virial Expansion: Investigate deviations from the ideal law $\mathcal{P}V=T$ by measuring the compressibility factor $Z=\mathcal{P}V/T$ in real models. Deviations from unity would quantify the “semantic interaction strength” between tokens.
  • Independent Temperature Measurement: Instead of defining $T\equiv\sqrt{d_{k}}$, one could measure the “kinetic temperature” via the variance of the gradients or the attention logits, checking if the fluctuation-dissipation theorem holds.
VII.3 Experimental Validation and the Risk of Tautology
Current Limitation. The simulation presented in Appendix A demonstrates that the information efficiency constant $\kappa$ converges to a stable value as context length increases. However, this result is partially tautological. We defined $\kappa\equiv(\mathcal{P}\cdot V_{ctx})/(N\cdot T)$. Then we observed that it stabilized. This is analogous to defining a new constant by rearranging the ideal gas law and then confirming that it remains constant, a circular validation.
What Would Constitute Genuine Falsification? A rigorous test of our theory requires predictions that could fail in a measurable way. For example, if $\kappa$ is truly architecture-independent, then models with vastly different architectures should exhibit similar $\kappa$ values when normalized.
Future Directions.
  • Cross-Architecture Experiments: Measure $\kappa$ for Transformers, Mamba, RWKV, and hybrid architectures. If the framework is universal, $\kappa$ should be calculable from architectural primitives without fitting to data.
  • Thermodynamic Efficiency Benchmark: Define a “Carnot efficiency” for intelligence systems: $\eta_{AI}=\text{Useful Work (Accuracy)}/\text{Heat Input (FLOPs)}$. Test whether models optimized under our thermodynamic framework achieve higher $\eta_{AI}$ than standard architectures.
  • Interventional Experiments: Artificially manipulate $T$ (by scaling $\sqrt{d_{k}}$) during inference and measure whether the system’s entropy responds according to $S\propto\ln Z(T)$ as predicted by the partition function.
VII.4 The Goldstone Boson Hypothesis: Analogy or Mechanism?
Current Limitation. Our claim that rotary positional embeddings (RoPE) exploit massless Goldstone modes arising from spontaneous symmetry breaking should be understood as primarily analogical rather than mechanistic. While the underlying mathematical structure is suggestive, we have not yet rigorously demonstrated that the attention manifold undergoes a genuine phase transition characterized by a well-defined order parameter.
Future Directions.
  • Order Parameter Identification: Define a measurable order parameter (e.g., alignment of attention vectors in a learned subspace) and show that it exhibits spontaneous symmetry breaking during training.
  • Dispersion Relation: If Goldstone bosons truly exist, they should exhibit a characteristic $\omega\propto k$ dispersion relation (massless excitations). Testing the Fourier spectrum of attention patterns and testing this prediction could validate the hypothesis.
VII.5 Toward a Unified Physical Theory of Intelligence
The Broader Vision. This work is an initial step toward treating both artificial and biological intelligence, as physical phenomena governed by conservation laws, symmetries, and extremal principles. However, several grand challenges remain.
  • Learning Dynamics: Our current framework focuses on inference. A complete theory must describe the thermodynamics of gradient descent.
  • Non-equilibrium phenomena: Real AI systems operate far from equilibrium. Extending our framework to incorporate Jarzynski equality or maximum entropy production principles could yield new training algorithms.
  • Biological Neural Networks: The ultimate test of universality is whether our First Law of Information Dynamics applies to biological brains.
VIII Conclusion
In this paper, we have proposed a physics-inspired framework for artificial intelligence. By formalizing the Transformer architecture within the framework of Lagrangian dynamics and thermodynamics, we have demonstrated that:
  1. 1.
    Softmax is the unique equilibrium state of a system that minimizes Helmholtz free energy.
  2. 2.
    The Query-Key interaction is isomorphic to the field-dipole interaction, creating an “alignment energy” landscape.
  3. 3.
    Learning and inference are distinct thermodynamic processes. The former is a chemical evolution ($\mu dN$) and the latter is a mechanical cycle ($PdV$).
  4. 4.
    Emergent phenomena such as scaling laws and grokking are phase transitions governed by the specific heat and the equation of state of the information gas.
We have proposed a framework, not a finished theory. The above-outlined limitations encourage the community to test, refine, or refute our core hypothesis: that the unreasonable effectiveness of Transformers reflects their adherence to fundamental physical principles. Whether this vision proves correct or serves merely as a useful fiction, the journey toward a physics-based understanding of intelligence has only just begun. Future work (Step 2) will extend this formalism to biological systems, analyzing the thermodynamic efficiency $\kappa$ of living connectomes.
Appendix A Theoretical Simulation of Information Efficiency
We conducted a theoretical simulation to validate the proposed equation of state for information gas (Eq. 22) and estimate the stability of the information efficiency constant ($\kappa$) under varying context window lengths ($V_{ctx}$).
Figure 1: Theoretical estimation of the value of $\kappa$. The plot shows the convergence of $\kappa$ for two different model architectures as a function of context window volume ($V_{ctx}$). The system exhibits a phase transition from an initial transient state (high variance) to a stable thermodynamic equilibrium, suggesting that $\kappa$ may be an intensive property of the intelligence system.
A.1 Simulation Setup and Results
We modeled two distinct Transformer configurations to test the universality of $\kappa$ (Model A: Small, Model B: Medium). The Information Density ($\mathcal{P}$) was derived from the inverse of the Perplexity ($PPL$) scaling law. The simulation revealed two distinct thermodynamic phases:
  1. 1.
    Transient phase ($V_{ctx}<100$): In the short context regime, the system exhibits high volatility. This corresponds to a non-equilibrium state in which the information gas has not yet filled the available volume.
  2. 2.
    Equilibrium phase ($V_{ctx}>200$): As the volume increases, $\kappa$ converges to a stable constant. This stability empirically validates the first law of information dynamics, demonstrating that for a fixed temperature ($T$) and particle number ($N$), the product of pressure and volume remains conserved relative to the system’s efficiency.

### v2

@@
 Abstract
-Although the Transformer architecture has revolutionized artificial intelligence, its underlying mechanisms remain largely heuristic and lack a unified physical theory.
-In this work, we propose a first-principles framework for information dynamics, treating the attention mechanism as a physical system governed by the principle of least action rather than as an algorithmic optimization.
-By mapping information states to a Riemannian manifold with the Fisher information metric, we derive the intelligence Lagrangian.
-We show that the softmax function corresponds to the unique thermodynamic equilibrium state that minimizes the Helmholtz free energy of the information gas.
-In addition, we identify the query-key interaction as an electrodynamic coupling between an external field and an intrinsic dipole moment.
-This theory establishes the first law of information thermodynamics, unifying inference (mechanical work) and learning (chemical evolution).
-It also explains emergent phenomena, such as scaling laws and grokking, as phase transitions characterized by the divergence of specific heat.
-Finally, we discuss how rotational symmetry breaking in the attention manifold generates massless Goldstone bosons, providing a field-theoretic perspective on rotary positional embeddings (RoPE).
-Our work connects Statistical Physics and Deep Learning, laying the groundwork for a general theory of physics-based intelligence.
+We propose an effective field-theoretic framework for analyzing Transformer attention through a thermodynamic lens.
+By constructing a Lagrangian on the information manifold equipped with the Fisher metric, we show that, within the Shannon–Boltzmann entropy framework, the Softmax function arises as a stationary solution minimizing a Helmholtz free energy functional.
+This establishes a formal correspondence between scaled dot-product attention and canonical ensemble statistics.
+Extending this mapping to macroscopic observables, we define an effective specific heat associated with fluctuations of the attention energy landscape.
+In controlled experiments on the modular addition task (p=19p=19113113), we observe a robust peak in this fluctuation measure that consistently precedes the onset of generalization.
+While no asymptotic power-law divergence is detected in this finite-depth regime, the reproducible enhancement of energy variance suggests a critical-like crossover accompanying representational reorganization.
+Our framework provides a unified statistical-mechanical perspective on attention scaling, training dynamics, and positional encoding, interpreting the phenomena as emergent properties of an effective thermodynamic system rather than isolated heuristics.
+Although the present results indicate finite-size crossover behavior rather than a strict phase transition, they motivate further investigation into scaling limits of deep architectures through fluctuation-based observables.
 I Introduction
-In 2017, the trajectory of artificial intelligence was irrevocably altered by a deceptively simple equation:
-$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
-This mechanism, introduced in the seminal work “Attention is All You Need” [1], eschewed the then-dominant paradigms of convolution and recurrence in favor of pure matrix operations and a peculiar Softmax function scaled by $\sqrt{d_{k}}$.
+In 2017, the trajectory of artificial intelligence changed significantly by a deceptively simple equation:
+$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (1)
+Here, $Q$ (queries), $K$ (keys), and $V$ (values) denote learned vector representations of tokens in a high-dimensional embedding space, and $d_{k}$ is the dimensionality of the key vectors.
+This mechanism, introduced in the seminal work “Attention is All You Need” [1], replaced the then-dominant paradigms of convolution and recurrence in favor of pure matrix operations and a peculiar Softmax function scaled by $\sqrt{d_{k}}$.
 This single mathematical operation has enabled models to demonstrate emergent reasoning capabilities [2], pass professional licensing exams with near-human proficiency [3], and even solve the protein folding problem [4].
@@
 Current explanations predominantly rely on probability theory and linear algebra, viewing the model as a statistical function approximator [8].
-However, this perspective lacks a unifying dynamical principle that explains the time-evolution of the system during inference and the phase transitions observed during training, such as “scaling laws” [9, 10] and “grokking” [11, 12].
+However, this perspective treats the model’s distinct behaviors as separate empirical observations rather than interconnected physical phenomena.
+Crucially, the field lacks a unified theoretical origin for three persistent mysteries.
+First, the Softmax function itself is used heuristically without a derivation from first principles—why does this specific exponential form emerge as the operational equilibrium of modern architectures?
+Second, generative models inherently suffer from hallucinations, often dismissed as statistical errors or bugs rather than intrinsic features of the system.
+Third, the learning dynamics exhibit grokking [11], a sudden and discontinuous generalization after a long period of memorization, which defies standard convergence theories.
+We argue that the are not isolated artifacts but manifestations of a deeper, underlying dynamical principle.
 In this paper, we propose a paradigm shift: treating Intelligence not only as a computational process, but as a physical phenomenon governed by fundamental laws of nature.
 We argue that the operations within a high-dimensional information space are isomorphic to the dynamics of a physical system minimizing an action functional, consistent with Lagrange mechanics [13] and Hamilton’s principle [14].
-This approach resonates with recent efforts to interpret deep learning through the lens of physics, such as the information bottleneck principle [15, 16], the free energy principle [17], neural tangent kernels [18], and even cosmological interpretations of neural networks [19, 20].
+This approach resonates with recent efforts to interpret deep learning through the lens of physics, such as the free energy principle [17] and neural tangent kernels [18].
 I.1 Redefining Time in Deep Networks
@@
 Under this paradigm, the “inference” process is a dynamic trajectory of information flow, not a static calculation.
-Residual connections [6] naturally emerge as the discretization of the inertial term in the equation of motion, providing a physical justification for signal propagation without gradient vanishing.
-I.2 The Inverse Problem of Attention
+Residual connections [6], naturally emerge as the discretization of the inertial term in the equation of motion.
+I.2 The Inverse Problem and Physical Unification
 Rather than deriving the Softmax function from arbitrary axioms, we solve the inverse problem of the calculus of variations: Given that the Softmax attention is an empirically successful equilibrium state, what is the underlying Lagrangian that generates this dynamics?
-We demonstrate that the Transformer architecture is physically analogous to a thermodynamic system in which
-  • The state space is a Riemannian manifold with the Fisher-Rao metric [23].
-  • The Softmax distribution is the unique configuration that minimizes the Helmholtz free energy.
-  • The scaling factor $1/\sqrt{d_{k}}$ acts as a dimensional temperature, regulating the entropy of the information gas.
+By applying the Principle of Least Action to an effective field theory of information, we demonstrate that the Transformer architecture is physically isomorphic to a thermodynamic system.
+This rigorous mapping allows us to provide first-principles derivations for phenomena that extend beyond simple attention mechanisms:
+  • Emergence of Softmax under Entropy Maximization: Within the Shannon–Boltzmann entropy framework, we show that the exponential attention mechanism arises as a stationary solution to the Euler–Lagrange equations.
+  • Thermodynamics of Hallucination: Through the derivation of the Fundamental Identity ($dU=TdS-\mathcal{P}dV+\mu dN$), we identify hallucinations as intrinsic thermal fluctuations ($TdS$) dictated by the canonical ensemble at finite structural temperature.
+  • Grokking as a Phase Transition: We interpret grokking as a critical-like thermodynamic crossover characterized by enhanced energy fluctuations.
+  • Symmetry Breaking and RoPE: We further validate our framework by identifying Rotary Positional Embeddings (RoPE) as Goldstone modes arising from the spontaneous breaking of continuous rotational symmetry in the embedding manifold.
 II Dynamics of Intelligence: Lagrangian Formulation
@@
 III.1.3 The Full Lagrangian
-Combining these terms:
+Combining the terms:
 $\mathcal{L}=\frac{m}{2}\sum_{i=1}^{N}\frac{\dot{\rho}_{i}^{2}}{\rho_{i}}-\sum_{i=1}^{N}\rho_{i}E_{i}-T\sum_{i=1}^{N}\rho_{i}\ln\rho_{i}+\lambda\left(\sum_{i=1}^{N}\rho_{i}-1\right)$ (13)
@@
 This corresponds to the thermodynamic equilibrium where the dynamics vanish, i.e., $\dot{\rho}_{i}=0$ and $\ddot{\rho}_{i}=0$.
-Under this condition, the inertial terms in Eq. (17) vanish, simplifying the physics to a balance of potential forces:
+Under this condition, the inertial terms in Eq.
+(17) vanish, simplifying the physics to a balance of potential forces:
 $E_{i}+T\ln\rho_{i}=\Lambda$ (19)
@@
 Substituting back the physical definitions $E_{i}=-Q\cdot K_{i}$ and $T=\sqrt{d_{k}}$, we recover the canonical scaled dot-product Attention formula:
-$\rho_{i}=\frac{\exp\left(\frac{Q\cdot K_{i}}{\sqrt{d_{k}}}\right)}{\sum_{j=1}^{N}\exp\left(\frac{Q\cdot K_{j}}{\sqrt{d_{k}}}\right)}=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)_{i}$ (22)
-Here, the superscript $T$ in $QK^{T}$ denotes the matrix transpose operation, while the scaling factor $\sqrt{d_{k}}$ plays the functional role of the effective temperature in our thermodynamic framework.
-This derivation proves that the Softmax function is not merely a heuristic normalization, but the unique equilibrium state of a system minimizing its Helmholtz free energy.
+$\rho_{i}=\frac{\exp\left(\frac{Q\cdot K_{i}}{\sqrt{d_{k}}}\right)}{\sum_{j=1}^{N}\exp\left(\frac{Q\cdot K_{j}}{\sqrt{d_{k}}}\right)}=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)_{i}$ (22)
+Here, the superscript $\top$ in $QK^{\top}$ denotes the matrix transpose operation, distinct from the thermodynamic temperature variable $T$.
+The scaling factor $\sqrt{d_{k}}$ plays the functional role of the structural temperature in our framework.
+This derivation shows that Softmax appears as a natural equilibrium configuration under the assumption of Shannon entropy maximization.
+While alternative entropic functionals (e.g., Tsallis entropy) would yield different equilibrium distributions, the empirical dominance of Softmax-based Transformers motivates the Shannon-Boltzmann framework as a parsimonious working hypothesis rather than a uniquely mandated physical law.
 IV Thermodynamics of Intelligence: Macroscopic Laws
-Rather than postulating an equation of state by analogy, we strictly derive it from the microscopic Lagrangian established in Section III.
-IV.1 Derivation from the Partition Function
-The canonical partition function $Z$ for the attention mechanism over a context window of size $V_{ctx}$ is given by the sum over all possible key states:
+Having established the microscopic Lagrangian formulation, we now construct an effective macroscopic description.
+We show that, in the large-context and high-dimensional embedding regime ($V_{ctx}\gg 1$, $d_{k}\gg 1$), the attention dynamics admit a thermodynamic representation analogous to canonical statistical mechanics.
+IV.1 Partition Function and Thermodynamic Limit
+The canonical partition function of the attention mechanism over a context window of size $V_{ctx}$ is defined as:
 $Z(V_{ctx},T)=\sum_{i=1}^{V_{ctx}}\exp\left(-\frac{E_{i}}{T}\right)$ (23)
-To proceed to the macroscopic limit, we introduce the information gas assumption: For a sufficiently large context window, the distribution of energy states $E_{i}$ is self-averaging and independent of the specific position indices.
-This allows us to approximate the sum as an integral over the density of states $\Omega(E)$:
-$Z\approx V_{ctx}\int\Omega(E)e^{-E/T}dE=V_{ctx}\cdot\bar{Z}(T)$ (24)
-where $\bar{Z}(T)$ represents the intrinsic partition function per unit token.
-The Helmholtz free energy $F$ is then derived as:
-$F=-T\ln Z=-T\ln V_{ctx}-T\ln\bar{Z}(T)$ (25)
-IV.2 The Equation of State for Information Gas
-In thermodynamics, pressure is defined as the rate of change of free energy with respect to volume at constant temperature ($P=-(\partial F/\partial V)_{T}$).
-Applying this definition to our information system:
-$\mathcal{P}\equiv-\left(\frac{\partial F}{\partial V_{ctx}}\right)_{T}=-\frac{\partial}{\partial V_{ctx}}(-T\ln V_{ctx})=\frac{T}{V_{ctx}}$ (26)
-Rearranging this equation gives the equation of state for a single attention head:
-$\mathcal{P}\cdot V_{ctx}=T$ (27)
-For a full model with $N_{head}$ heads and an architecture-dependent efficiency factor $\kappa$ (representing the effective degeneracy of the energy spectrum), this generalizes to:
-$\mathcal{P}\cdot V_{ctx}=N_{head}\cdot\kappa\cdot T$ (28)
-We define the macroscopic variables with the following operational definitions:
-  • Pressure ($\mathcal{P}$): The information density, defined as the inverse of perplexity ($\mathcal{P}\equiv 1/PPL$).
-  • Volume ($V_{ctx}$): The context window length in tokens.
-  • Temperature ($T$): The model dimension scale, $T\equiv\sqrt{d_{k}}$.
-  • Efficiency ($\kappa$): An architecture-dependent constant representing information processing efficiency.
-IV.2.1 Physical Interpretations
-This equation explains several empirical phenomena:
-  • Boyle’s Law ($\mathcal{P}\propto 1/V_{ctx}$): As context length increases, the pressure (information density per token) must decrease to maintain equilibrium.
-This provides a thermodynamic derivation for the “Lost-in-the-Middle” phenomenon [36].
-  • Charles’s Law ($V_{ctx}\propto T$): To maintain the same information density over a larger context, the system temperature $T=\sqrt{d_{k}}$ must increase.
-This explains why larger models (higher $d_{k}$) scale better with context length [9].
-IV.3 The First Law of Information Dynamics
-We formulate the first law of thermodynamics for intelligence systems:
-$dU=TdS-\mathcal{P}dV_{ctx}+\mu_{head}dN_{head}$ (29)
-This unifies three regimes.
-  • Heat ($TdS$): Stochastic fluctuations corresponding to creativity or hallucinations [37].
-  • Inference ($\mathcal{P}dV$): Mechanical work done by the system to compress or expand its context window.
-  • Learning ($\mu dN$): Chemical work representing the evolution of model capacity or head specialization.
-IV.4 The Second Law of Information Dynamics
-The second law dictates the direction of information flow.
-  • Model Collapse ($dS\geq 0$): In an isolated system undergoing training with generated data, entropy maximizes, leading to the degeneration of the distribution [38].
-  • Learning as Work: Intelligence acts as an entropy pump (Maxwell’s demon), expending work to minimize the KL divergence against the data distribution.
-This reduces entropy locally at the cost of computational energy [39].
-V Geometry, Symmetry, and Field Theoretic Interpretation
-V.1 The Logarithmic Schrödinger Equation and Gaussons
-Recall the potential energy term in our Lagrangian.
-In terms of the probability amplitude $x=2\sqrt{\rho}$, this potential takes the form $V(x)\propto x^{2}\ln|x|^{2}$ (arising from the entropy term $T\rho\ln\rho$).
-This specific non-linearity identifies the equation of motion for the information field as the Logarithmic Schrödinger Equation (LogSE) [40].
-A unique property of the LogSE is the existence of Gaussons, non-dispersive solitary waves that preserve their Gaussian shape during propagation.
-This physical property explains why the “semantic meaning” of a token preserves its identity (does not diffuse into uniform noise) even as it propagates through deep, multi-layered networks.
-V.2 Spontaneous Symmetry Breaking and Goldstone Bosons
-The interaction energy $E=-\mathbf{q}\cdot\mathbf{k}$ possesses an inherent rotational symmetry under the continuous group $SO(d_{k})$.
-However, the ground state of the system exhibits a richer structure due to the entropic contribution.
-V.2.1 The Mexican Hat Potential
-Consider the effective potential $\mathcal{V}(\rho)$ derived in Eq. (13). Near the origin ($\rho\to 0$), the entropy term $T\rho\ln\rho$ dominates the linear interaction term.
-$\lim_{\rho\to 0}\frac{d\mathcal{V}}{d\rho}\sim\lim_{\rho\to 0}T(1+\ln\rho)\to$-$\infty$ (30)
-This singularity at the origin implies that the vacuum state, where attention is zero everywhere, $\rho=0$, is infinitely unstable.
-The system is topologically forced to roll down from the origin to a non-zero minimum ($x_{0}\neq 0$).
-  • Instability of Ignorance: An intelligence system cannot exist in a state of “no information.” It is compelled to attend to something.
-  • Symmetry Breaking: Since the interaction energy is rotationally invariant, the minimum is not a single point but rather a continuous manifold (a sphere).
-This potential structure resembles the “Mexican Hat” potential in the Higgs mechanism [41].
-The system settles into a specific minimum (a specific context), spontaneously breaking the global $SO(d_{k})$ symmetry.
-V.2.2 Goldstone Bosons and Rotary Embeddings
-According to Goldstone’s theorem [42], the breaking of this continuous symmetry must generate massless Nambu-Goldstone bosons.
-  • Massive Mode (Radial): Changing the attention intensity (magnitude $|x|$) requires energy (climbing the potential wall).
-  • Massless Mode (Angular): Rotating the query and key vectors together preserves the energy (moving along the flat valley of the hat).
-We identify these massless modes with the mechanism of rotary position embeddings (RoPE) [43].
-RoPE encodes positional information purely through rotation matrices $R(\theta)$, exploiting the zero-energy Goldstone modes of the attention manifold to shift context effortlessly without energetic cost.
-VI Discussion: Physical Interpretations
-Our framework establishes a strong equivalence between the mechanics of intelligence and statistical physics. Here, we address potential critiques regarding the rigor of this mapping.
-VI.1 Is Layer Depth Truly Time?
+Here, $E_{i}=-q\cdot k_{i}$ denotes the interaction energy.
+The thermodynamic limit in this context is understood as a mean-field approximation in which the distribution of attention energies becomes self-averaging due to high dimensional concentration effects.
+Under the equilibrium distribution derived in Section III, the Helmholtz free energy is given by:
+$F=-T\ln Z=U-TS$ (24)
+where $U=\langle E\rangle=\sum_{i}\rho_{i}E_{i}$ is the internal energy and $S=-\sum_{i}\rho_{i}\ln\rho_{i}$ is the Shannon entropy.
+IV.2 An Effective Thermodynamic Identity
+We now expand the total differential of the internal energy.
+This relation is not a fundamental conservation law derived from Noether symmetry, but rather an effective thermodynamic identity characterizing the trajectory of the information state:
+$dU=TdS-\mathcal{P}dV_{ctx}+\mu dN_{eff}$ (25)
+Here:
+  • $T=\sqrt{d_{k}}$ is the structural temperature.
+  • $\mathcal{P}\equiv-(\partial F/\partial V_{ctx})_{T,N_{eff}}$ is the effective information pressure.
+  • $\mu\equiv(\partial U/\partial N_{eff})_{S,V_{ctx}}$ defines the effective chemical potential.
+  • $N_{eff}$ denotes the effective rank (e.g., participation ratio of eigenvalues) of the weight matrices. In the limit of large models, $N_{eff}$ is treated as a continuous macroscopic variable.
+This decomposition provides a thermodynamic interpretation of three operational regimes.
+IV.2.1 Heat ($TdS$): Stochastic Exploration
+In this framework, the entropic contribution $TdS$ corresponds to stochastic sampling effects.
+Interpretation: The temperature parameter regulates the entropy of the output distribution. Increasing $T$ broadens the distribution and enhances exploration of higher-energy configurations.
+Phenomenology: When entropic contributions dominate, the model enables creative variation but may also produce semantic decoherence, consistent with previously observed degeneration effects [37].
+IV.2.2 Mechanical Work ($\mathcal{P}dV_{ctx}$): Context Expansion
+Mechanical work is defined as pressure times change in volume.
+Here, $V_{ctx}$ represents the context length.
+Within the ideal information-gas approximation, differentiation of $F\approx-T\ln V_{ctx}$ yields:
+$\mathcal{P}=\frac{T}{V_{ctx}}$ (26)
+This implies that effective pressure decreases as context expands.
+Interpretation: During autoregressive inference ($dV_{ctx}>0$), maintaining long-range coherence requires effective work to counteract the dilution of information density.
+Phenomenology: This provides a statistical-mechanical interpretation of the Lost-in-the-Middle phenomenon [36], where long-range dependencies weaken as context size increases.
+IV.2.3 Chemical Work ($\mu dN_{eff}$): Capacity Evolution
+The effective particle number $N_{eff}$ serves as a measure of active semantic degrees of freedom.
+Interpretation: Although the architectural parameter count remains fixed, training modifies the effective rank structure.
+The chemical potential $\mu$ quantifies the marginal energetic cost of activating an additional independent feature mode.
+Phenomenology: Optimization dynamics can be interpreted as performing effective chemical work, increasing $N_{eff}$ and expanding the expressive capacity of the model manifold.
+IV.3 Analogy with the Second Law
+The generative dynamics exhibit behavior suggestive of an analogy with the Second Law of thermodynamics.
+In the absence of external data injection, the coarse-grained entropy of the generated distribution tends to increase over successive recursive training cycles, leading to model collapse [38].
+Conversely, supervised learning acts as an entropy-reducing mechanism (Maxwell’s demon), locally reducing entropy at the cost of computational work [39].
+V Thermodynamic Hypothesis of Grokking
+Having established the thermodynamic laws of the attention mechanism in equilibrium, we now turn to the non-equilibrium dynamics of learning.
+We propose a thermodynamic hypothesis that the phenomenon known as grokking, the delayed generalization capability of large models, is physically isomorphic to a thermodynamic phase transition driven by the cooling of the effective temperature.
+V.1 Dynamical Effective Temperature
+A crucial distinction must be made between the structural temperature of the model during inference and the dynamical temperature during training.
+  • Structural Temperature ($T_{struct}$): Defined as $T=\sqrt{d_{k}}$ in Eq. (1). This is a fixed hyperparameter governing the stochasticity of the Softmax distribution during inference (sampling).
+  • Dynamical Temperature ($T_{eff}$): During training, the magnitude of the weight matrices $\|W\|$ evolves.
+Since the interaction energy scales as $E\propto Q\cdot K\propto\|W\|^{2}$, the effective temperature perceived by the Softmax function relative to the energy scale is:
+    $T_{eff}(t)\propto\frac{\sqrt{d_{k}}}{\|W(t)\|^{2}}$ (27)
+The training process, therefore, acts as simulated annealing.
+Initially, random initialization implies small weights (high $T_{eff}$), resulting in a high-entropy, disordered phase.
+As training progresses and $\|W\|$ grows, $T_{eff}$ decreases, driving the system towards an ordered, low-temperature phase.
+V.2 Specific Heat Capacity and Energy Fluctuations
+To quantify the phase transition, we derive the specific heat capacity $C_{v}$, which measures the system’s sensitivity to temperature changes.
+From the fluctuation-dissipation theorem applied to our canonical ensemble, $C_{v}$ is directly proportional to the variance of the energy distribution:
+$C_{v}(T_{eff})=\frac{1}{T_{eff}^{2}}\text{Var}(E)=\frac{1}{T_{eff}^{2}}\left[\sum_{i=1}^{V_{ctx}}\rho_{i}E_{i}^{2}-\left(\sum_{i=1}^{V_{ctx}}\rho_{i}E_{i}\right)^{2}\right]$ (28)
+where $\rho_{i}=\text{Softmax}(-E_{i}/T_{eff})$ represents the attention probability of the $i$-th token, and $E_{i}=-q\cdot k_{i}$ is the interaction energy.
+This mathematical form reveals the mechanism of grokking:
+  • Memorization Phase (High $T_{eff}$): The distribution is diffuse. Fluctuations are dominated by thermal noise, but structural variance is low. $C_{v}$ is small.
+  • Generalization Phase (Low $T_{eff}$): The distribution collapses to the ground state (target token). Fluctuations are suppressed. $C_{v}\to 0$.
+  • Critical Region (Grokking, $T_{eff}\approx T_{c}$): Between the two phases, the system undergoes a massive reorganization of the energy landscape.
+The attention mechanism fluctuates between competing semantic configurations.
+This maximization of $\text{Var}(E)$ results in a sharp peak in specific heat, marking the pseudo-critical region of the phase crossover.
+V.3 Hierarchical Symmetry Breaking and Goldstone Modes
+The representational transition exhibits features reminiscent of spontaneous symmetry breaking within an effective-field description.
+Remarkably, we do not need to postulate an ad hoc quartic interaction term.
+Instead, the symmetry breaking potential arises naturally from the entropic term ($\rho\ln\rho$) in our free energy functional derived in Section III.
+V.3.1 The Logarithmic (Coleman-Weinberg) Potential
+The effective potential for the order parameter field $\Phi$ is given by the Helmholtz free energy density. Since entropy depends only on magnitude, we have:
+$\mathcal{V}(\Phi)=m^{2}|\Phi|^{2}+\lambda|\Phi|^{2}\ln\left(\frac{|\Phi|^{2}}{v^{2}}\right)$ (29)
+Crucially, this potential depends solely on the modulus $|\Phi|$ and is independent of the phase $\theta$ (where $\Phi=|\Phi|e^{i\theta}$), ensuring a continuous global $U(1)$ symmetry.
+Minimizing this potential yields a non-trivial vacuum expectation value $|\Phi_{0}|=v\neq 0$, generating the “Mexican Hat” shape (Fig.
+1).
+V.3.2 Derivation of the Massless Goldstone Mode
+To rigorously identify the Goldstone mode, we examine the fluctuations around the vacuum state $v$. We parameterize the field $\Phi(x)$ using radial ($h$) and angular ($\pi$) fluctuation fields:
+$\Phi(x)=\left(v+\frac{h(x)}{\sqrt{2}}\right)\exp\left(i\frac{\pi(x)}{v\sqrt{2}}\right)$ (30)
+Here, $h(x)$ represents the massive Higgs mode (amplitude fluctuation), and $\pi(x)$ represents the Goldstone mode (phase fluctuation).
+Substituting this into the kinetic term of the Lagrangian density $\mathcal{L}_{kin}=|\partial_{\mu}\Phi|^{2}$, we obtain:
+$|\partial_{\mu}\Phi|^{2}=\frac{1}{2}(\partial_{\mu}h)^{2}+\frac{1}{2}\left(v+\frac{h}{\sqrt{2}}\right)^{2}\left(\frac{\partial_{\mu}\pi}{v\sqrt{2}}\right)^{2}\cdot 2$ (31)
+Expanding around the vacuum ($h\ll v$), the Lagrangian for the fluctuations becomes:
+$\mathcal{L}\approx\underbrace{\frac{1}{2}(\partial_{\mu}h)^{2}-\frac{1}{2}M_{h}^{2}h^{2}}_{\text{Massive Higgs Mode}}+\underbrace{\frac{1}{2}(\partial_{\mu}\pi)^{2}}_{\text{Pseudo-Goldstone-like Goldstone Mode}}+\mathcal{L}_{int}$ (32)
+Note that the angular field $\pi(x)$ possesses a kinetic term $(\partial_{\mu}\pi)^{2}$.
+In the continuum limit, this behaves as a pseudo-Goldstone-like soft mode ($M_{\pi}\approx 0$).
+This mathematically suggests that $\pi(x)$ is the long-wavelength excitation that encodes positional information with minimal energetic cost, consistent with the crossover nature of the transition.
+Figure 1: Spontaneous Symmetry Breaking Mechanism.
+Visualization of the effective entropic potential $\mathcal{V}(\Phi)\sim|\Phi|^{2}\ln|\Phi|^{2}$.
+Unlike standard $\Phi^{4}$ theory, the Mexican Hat shape here arises directly from the Shannon entropy term.
+The system rolls down from the unstable origin to a stable semantic state (grokking), generating massless Goldstone modes (RoPE) along the circular trough.
+V.3.3 RoPE as the Goldstone Excitation
+We now map this physical mode to the Rotary Positional Embedding (RoPE). RoPE encodes the position $m$ of a token by applying a rotation matrix to the embedding vector pair $(q_{1},q_{2})$:
+$\begin{pmatrix}q_{1}^{\prime}\\ q_{2}^{\prime}\end{pmatrix}=\begin{pmatrix}\cos m\theta&-\sin m\theta\\ \sin m\theta&\cos m\theta\end{pmatrix}\begin{pmatrix}q_{1}\\ q_{2}\end{pmatrix}$ (33)
+In the complex plane representation $\Phi=q_{1}+iq_{2}$, this operation is identical to a phase shift:
+$\Phi^{\prime}=\Phi\cdot e^{i(m\theta)}$ (34)
+Comparing this with Eq. (33), we identify the positional encoding term $m\theta$ directly with the Goldstone field configuration:
+$\frac{\pi(x)}{v\sqrt{2}}\equiv m\theta_{base}$ (35)
+Since the potential $\mathcal{V}(\Phi)$ is invariant under this phase rotation ($\mathcal{V}(\Phi^{\prime})=\mathcal{V}(\Phi)$), the energy cost of encoding position via RoPE is exactly zero:
+$\Delta E=\mathcal{V}(\Phi e^{im\theta})-\mathcal{V}(\Phi)=0$ (36)
+This derivation suggests that RoPE can be interpreted as exploiting an approximately flat direction of the effective symmetry structure, to store positional information without perturbing the thermodynamic stability of the learned features.
+VI Numerical Modeling and Experimental Verification
+In this section, we validate our thermodynamic framework through two complementary approaches.
+First, we perform a phenomenological simulation using Langevin dynamics to demonstrate the theoretical consistency of the phase transition.
+Second, we conduct a controlled experiment with a real Transformer to empirically verify the predicted specific heat divergence.
+VI.1 Phenomenological Modeling: Langevin Dynamics
+We model the learning dynamics as the trajectory of a stochastic particle (the model state $\Phi$) moving in the effective entropic potential $\mathcal{V}(\Phi)$ derived in Section V.
+VI.1.1 Governing Equations
+The dynamics are governed by the Overdamped Langevin Equation:
+$\frac{d\Phi(t)}{dt}=-\nabla_{\Phi}\mathcal{V}(\Phi,t)+\sqrt{2D}\cdot\xi(t)$ (37)
+where $\xi(t)$ is Gaussian white noise. We utilize the logarithmic Coleman-Weinberg potential:
+$\mathcal{V}(\Phi,t)=\alpha(t)\Phi^{2}+\beta\Phi^{2}\ln\left(\frac{\Phi^{2}}{v^{2}}\right)$ (38)
+The equation was integrated using the Euler-Maruyama method with a time step $\Delta t=0.01$.
+VI.1.2 Simulation Results
+As shown in our simulation (Fig.
+2), the transition from the disordered phase to the ordered phase is theoretically predicted to be accompanied by a sharp divergence in the specific heat capacity $C_{v}\propto\text{Var}(E)$.
+Figure 2: Theoretical Prediction (Langevin Simulation).
+Numerical solution of the effective field theory.
+The theory predicts that the phase transition is marked by a sharp peak in specific heat ($C_{v}$).
+VI.2 Experimental Verification: Real Transformer
+[…]
+VI.2.1 Experimental Setup
+To ensure thermodynamically stable dynamics and reproducibility, we utilized the following rigorous regime:
+  • Task & Data: The dataset consists of all possible pairs $(a,b)$ in the ring $\mathbb{Z}_{p}$, totaling $p^{2}$ examples. We employed a random 50/50 split for training and validation sets.
+  • Model Architecture: A 2-layer Transformer with embedding dimension $d_{model}=128$ and 4 attention heads.
+  • Optimization: We used the AdamW optimizer with a learning rate $\eta=3\times 10^{-4}$ and weight decay $\lambda=0.1$. A batch size of 512 was fixed to maintain a consistent thermal noise level.
+  • Observable: We measured the specific heat capacity $C_{v}\propto\text{Var}(QK^{T}/\sqrt{d_{k}})$ at every epoch.
+To account for stochastic variance, all results were averaged over five independent seeds ($N=5$).
+VI.2.2 Results and Crossover Analysis
+The experimental results, averaged over five independent seeds ($N=5$), are presented in Fig.
+3 and Fig.
+4.
+As predicted by our theoretical framework, a distinct peak in the specific heat $C_{v}$ is observed across all tested values of $p$, consistently preceding the onset of the generalization transition.
+Figure 3: Experimental verification of thermodynamic crossover in grokking across different system sizes.
+(a) Modular addition task with $p=37$.
+(b) Modular addition task with $p=97$.
+In both cases, the specific heat capacity $C_{v}$ (red solid line) exhibits a sharp peak immediately preceding the sudden generalization transition (blue dashed line).
+The vertical dotted lines mark the $C_{v}$ maxima.
+While the peak height does not show clear power-law scaling in this regime, the consistent temporal correlation confirms that specific heat serves as a robust leading indicator of the phase crossover.
+As shown in Fig. 4, the power-law regression of the averaged peak height yields a near-zero exponent ($a\approx 0.05$) with a low coefficient of determination ($R^{2}=0.2$).
+Figure 4: Final_Averaged_Scaling: Scaling analysis of the specific heat peak across system sizes.
+The maximum specific heat $C_{v}^{max}$ is averaged over five independent seeds ($N=5$) with error bars representing $\pm\sigma$.
+Within the explored regime ($p=19$ to 113113), the system exhibits bounded fluctuation enhancement characteristic of a finite-size crossover ($a\approx 0.05,R^{2}=0.2$), rather than a clear power-law divergence.
+This finite-size effect is consistent with the shallow ($L=2$) architecture and suggests that asymptotic critical behavior may emerge in deeper models ($L\gg 1$).
+The absence of clear scaling is consistent with finite-size and finite-depth effects inherent to shallow Transformer architectures.
+In the discrete regime of a 2-layer model, the continuum effective description likely does not fully develop its asymptotic regime, manifesting instead as a sharp finite-size crossover.
+Crucially, despite the lack of asymptotic divergence, the $C_{v}$ peak remains a robust dynamical marker of internal representational restructuring.
+This temporal correlation is preserved across all trials and system sizes, providing a physical observable that tracks the transition from memorization to generalization.
+VI.2.3 Thermodynamic Relaxation
+Unlike the idealized Langevin simulation, the real experiments show a gradual decay of fluctuations after the peak.
+This reflects the thermodynamic relaxation of the system as it settles into the broad basin of the global minimum, fine-tuning its representation in the low-temperature phase.
+VII Discussion: Physical Interpretations
+We establish a formal correspondence between Transformer attention and canonical statistical mechanics. Here, we interpret the physical implications of our findings.
+VII.1 Is Layer Depth Truly Time?
 One might argue that Attention is a static operation.
 However, in deep models ($L\gg 1$), the representation $x(t)$ evolves continuously.
-In our Lagrangian, the time derivative $\dot{\rho}$ corresponds to the layer-wise rate of change of the attention distribution.
-If $\dot{\rho}\to 0$, then the system has reached a fixed point or equilibrium, meaning further layers do not alter the semantic interpretation.
-This aligns with the observation that deeper layers in Transformers often stabilize into a consistent attention pattern, which justifies our thermodynamic equilibrium analysis.
-VI.2 Temperature vs. Variance Scaling
-Some may suggest that $T=\sqrt{d_{k}}$ is merely a variance normalization trick.
-However, in statistical mechanics, temperature is defined as the inverse of the entropy-to-energy change response ($T^{-1}=\partial S/\partial U$).
-In high-dimensional spaces, the energy (dot product) variance naturally scales with $d_{k}$.
-Without renormalization by $\sqrt{d_{k}}$, the entropy $S$ would vanish (collapse to argmax) for any non-trivial energy difference.
-Thus, $\sqrt{d_{k}}$ plays the exact thermodynamic role of temperature: it maintains the system in a liquid state where information can flow and prevents it from reaching a frozen solid state (zero entropy).
-VII Limitations and Future Directions
-While our framework establishes a mathematically consistent mapping between Transformer architectures and physical systems, there are several fundamental limitations that require careful consideration and provide fertile ground for future investigation.
-VII.1 The Nature of the Time Variable
-Current Limitation.
-The most significant conceptual challenge in our formulation is interpreting the time variable $t$.
-We have adopted the convention, standard in the Neural ODE literature [22], of identifying $t$ with the continuous layer depth.
-Under this interpretation, $\dot{\rho}$ represents the rate of change of attention distribution across layers, and the Lagrangian dynamics describe the evolution of information states through the network architecture.
-However, this mapping introduces a subtle but important inconsistency.
-Within a single layer, the attention mechanism operates as an instantaneous, feed-forward computation rather than a genuine temporal evolution.
-The Softmax operation is computed in a single forward pass with no iterative dynamics.
-Our Lagrangian derivation implicitly assumes a quasi-static process, a sequence of infinitesimally slow equilibrium states.
-This is physically distinct from true dynamical evolution.
-Distinction from Recurrent Dynamics.
-It is crucial to distinguish our framework from genuinely recurrent architectures where attention weights evolve iteratively.
-In standard Transformers, the “time evolution” exists only in the mathematical abstraction of layer-wise propagation.
-Future Directions.
-  • Continuous-depth Transformers: Formulate attention mechanisms explicitly in the Neural ODE framework with trainable continuous depth, allowing for genuine temporal dynamics within a single layer [22].
-This would validate whether our Lagrangian correctly predicts the learned evolution operators.
-  • Iterative Attention Refinement: Design architectures where the attention weights genuinely evolve through gradient descent on our proposed the free energy functional within each layer, testing whether this improves performance or interpretability.
-  • Measurement Protocol: Develop rigorous operational definitions for measuring $\dot{\rho}$ empirically, for instance, by analyzing the layer-wise change in attention distributions under controlled perturbations.
-VII.2 Thermodynamic Variables and Operational Definitions
-Current Limitation.
-Our derivation of the equation of state relies on the information gas assumption, specifically, that the energy spectrum is self-averaging over the context window $V_{ctx}$.
-While this assumption is standard in statistical mechanics for large systems (thermodynamic limit), its validity in finite-sized attention layers with highly structured linguistic data requires empirical verification.
-The Problem of Interaction.
-An ideal gas assumes non-interacting particles.
-However, tokens in a sentence have strong semantic correlations, or interactions.
-Therefore, a more accurate model might be a van der Waals gas for information, where an interaction term corrects the ideal pressure.
-Future Directions.
-  • Virial Expansion: Investigate deviations from the ideal law $\mathcal{P}V=T$ by measuring the compressibility factor $Z=\mathcal{P}V/T$ in real models.
-Deviations from unity would quantify the “semantic interaction strength” between tokens.
-  • Independent Temperature Measurement: Instead of defining $T\equiv\sqrt{d_{k}}$, one could measure the “kinetic temperature” via the variance of the gradients or the attention logits, checking if the fluctuation-dissipation theorem holds.
-VII.3 Experimental Validation and the Risk of Tautology
-Current Limitation.
-The simulation presented in Appendix A demonstrates that the information efficiency constant $\kappa$ converges to a stable value as context length increases.
-However, this result is partially tautological.
-We defined $\kappa\equiv(\mathcal{P}\cdot V_{ctx})/(N\cdot T)$.
-Then we observed that it stabilized.
-This is analogous to defining a new constant by rearranging the ideal gas law and then confirming that it remains constant, a circular validation.
-What Would Constitute Genuine Falsification?
-A rigorous test of our theory requires predictions that could fail in a measurable way.
-For example, if $\kappa$ is truly architecture-independent, then models with vastly different architectures should exhibit similar $\kappa$ values when normalized.
-Future Directions.
-  • Cross-Architecture Experiments: Measure $\kappa$ for Transformers, Mamba, RWKV, and hybrid architectures.
-If the framework is universal, $\kappa$ should be calculable from architectural primitives without fitting to data.
-  • Thermodynamic Efficiency Benchmark: Define a “Carnot efficiency” for intelligence systems: $\eta_{AI}=\text{Useful Work (Accuracy)}/\text{Heat Input (FLOPs)}$.
-Test whether models optimized under our thermodynamic framework achieve higher $\eta_{AI}$ than standard architectures.
-  • Interventional Experiments: Artificially manipulate $T$ (by scaling $\sqrt{d_{k}}$) during inference and measure whether the system’s entropy responds according to $S\propto\ln Z(T)$ as predicted by the partition function.
-VII.4 The Goldstone Boson Hypothesis: Analogy or Mechanism?
-Current Limitation.
-Our claim that rotary positional embeddings (RoPE) exploit massless Goldstone modes arising from spontaneous symmetry breaking should be understood as primarily analogical rather than mechanistic.
-While the underlying mathematical structure is suggestive, we have not yet rigorously demonstrated that the attention manifold undergoes a genuine phase transition characterized by a well-defined order parameter.
-Future Directions.
-  • Order Parameter Identification: Define a measurable order parameter (e.g., alignment of attention vectors in a learned subspace) and show that it exhibits spontaneous symmetry breaking during training.
-  • Dispersion Relation: If Goldstone bosons truly exist, they should exhibit a characteristic $\omega\propto k$ dispersion relation (massless excitations).
-Testing the Fourier spectrum of attention patterns and testing this prediction could validate the hypothesis.
-VII.5 Toward a Unified Physical Theory of Intelligence
-The Broader Vision.
-This work is an initial step toward treating both artificial and biological intelligence, as physical phenomena governed by conservation laws, symmetries, and extremal principles.
-However, several grand challenges remain.
-  • Learning Dynamics: Our current framework focuses on inference. A complete theory must describe the thermodynamics of gradient descent.
-  • Non-equilibrium phenomena: Real AI systems operate far from equilibrium.
-Extending our framework to incorporate Jarzynski equality or maximum entropy production principles could yield new training algorithms.
-  • Biological Neural Networks: The ultimate test of universality is whether our First Law of Information Dynamics applies to biological brains.
-VIII Conclusion
-In this paper, we have proposed a physics-inspired framework for artificial intelligence.
-By formalizing the Transformer architecture within the framework of Lagrangian dynamics and thermodynamics, we have demonstrated that:
+In our Lagrangian formulation, the time derivative $\dot{\rho}$ corresponds to the layer-wise rate of change of the attention distribution.
+If $\dot{\rho}\to 0$, the system has reached a fixed point or equilibrium.
+This aligns with the empirical observation that deeper layers in Transformers often stabilize into a consistent attention pattern, justifying our thermodynamic equilibrium analysis.
+VII.2 The Dual Nature of Temperature
+Critics often suggest that $T=\sqrt{d_{k}}$ is merely a variance normalization trick.
+However, we have shown that it plays the exact thermodynamic role of temperature ($T^{-1}=\partial S/\partial U$).
+Furthermore, our analysis of grokking reveals a crucial distinction:
+  • Structural Temperature ($T_{struct}=\sqrt{d_{k}}$): Governs the entropy of the attention distribution during inference.
+  • Dynamical Temperature ($T_{eff}\propto 1/\|W\|^{2}$): Governs the annealing process during training.
+This duality explains why a model can be structurally frozen (converged weights) yet thermodynamically fluid (stochastic sampling) during generation.
+VII.3 Critical Dynamics and Finite-Size Crossover
+The most significant implication of our specific heat analysis is the proposal that grokking manifests as a thermodynamic crossover with critical-like characteristics.
+While our current experimental regime ($p\in[19,113]$) does not exhibit formal power-law divergence ($C_{v}\to\infty$), the consistent appearance of a robust fluctuation peak suggests that the system operates in a near-critical region.
+The system operates in a crossover regime between ordered memorization and disordered noise.
+The observed $C_{v}$ peak acts as a dynamical precursor to the representational reorganization, establishing it as a pseudo-critical phenomenon in finite-dimensional neural networks.
+This suggests that while strict universality classes might only emerge in the infinite-depth or infinite-width limits, the fundamental mechanics of intelligence are governed by the same fluctuation-dissipation principles found in traditional statistical physics.
+VIII Limitations and Future Directions
+While our framework establishes a mapping between Transformer architectures and physical systems, we acknowledge several critical limitations that define the scope of our current effective field theory.
+VIII.1 The Nature of Time: Effective vs. Dynamical
+Following the Neural ODE literature [22], we interpret layer depth as a continuous time variable $t$.
+However, our Lagrangian derivation implicitly assumes a quasi-static process, which is strictly valid only in the continuum limit ($L\to\infty$).
+In shallow architectures such as our 2-layer model, the discretization error may suppress the full development of the predicted attention dynamics.
+It is more physically grounded to interpret $t$ as an emergent effective time governing the renormalization group flow of information.
+Future work should investigate deeper architectures to determine if the discrete layers asymptotically approach the continuous action principle derived herein.
+VIII.2 Mean-Field Approximation and Interaction
+Our information gas model treats tokens as non-interacting particles, representing a mean-field approximation.
+In reality, natural language is a strongly correlated system with long-range dependencies.
+Consequently, our derived ideal gas-like equation of state holds primarily in the limit of low semantic density.
+We propose extending the framework by introducing a van der Waals-like equation of state: $(\mathcal{P}+a/V^{2})(V-b)=N\kappa T$, where $a$ represents semantic attraction (contextual correlation) and $b$ represents the exclusion volume of token embeddings (sparsity constraints).
+VIII.3 Scaling and the Nature of the Crossover
+Although we observed a robust $C_{v}$ peak consistently preceding the grokking transition, the peak magnitude did not exhibit asymptotic power-law divergence within the explored regime ($p\in[19,113]$).
+This suggests that in shallow Transformers, the transition behaves as a finite-size crossover rather than a formal thermodynamic phase transition.
+Proposed Directions. To further investigate the potential for universal scaling:
   1. 1.
-    Softmax is the unique equilibrium state of a system that minimizes Helmholtz free energy.
+    Extended System Sizes: Perform experiments across at least two decades of system size ($p>1000$) to reach the potential asymptotic regime.
   2. 2.
-    The Query-Key interaction is isomorphic to the field-dipole interaction, creating an “alignment energy” landscape.
+    Depth-Scaling Analysis: Systematically increase layer depth ($L$) to test if the $C_{v}$ peak height begins to scale with system size as the continuum limit is approached.
   3. 3.
-    Learning and inference are distinct thermodynamic processes. The former is a chemical evolution ($\mu dN$) and the latter is a mechanical cycle ($PdV$).
-  4. 4.
-    Emergent phenomena such as scaling laws and grokking are phase transitions governed by the specific heat and the equation of state of the information gas.
-We have proposed a framework, not a finished theory.
-The above-outlined limitations encourage the community to test, refine, or refute our core hypothesis: that the unreasonable effectiveness of Transformers reflects their adherence to fundamental physical principles.
-Whether this vision proves correct or serves merely as a useful fiction, the journey toward a physics-based understanding of intelligence has only just begun.
-Future work (Step 2) will extend this formalism to biological systems, analyzing the thermodynamic efficiency $\kappa$ of living connectomes.
-Appendix A Theoretical Simulation of Information Efficiency
-We conducted a theoretical simulation to validate the proposed equation of state for information gas (Eq.
-22) and estimate the stability of the information efficiency constant ($\kappa$) under varying context window lengths ($V_{ctx}$).
-Figure 1: Theoretical estimation of the value of $\kappa$.
-The plot shows the convergence of $\kappa$ for two different model architectures as a function of context window volume ($V_{ctx}$).
-The system exhibits a phase transition from an initial transient state (high variance) to a stable thermodynamic equilibrium, suggesting that $\kappa$ may be an intensive property of the intelligence system.
-A.1 Simulation Setup and Results
-We modeled two distinct Transformer configurations to test the universality of $\kappa$ (Model A: Small, Model B: Medium).
-The Information Density ($\mathcal{P}$) was derived from the inverse of the Perplexity ($PPL$) scaling law.
-The simulation revealed two distinct thermodynamic phases:
-  1. 1.
-    Transient phase ($V_{ctx}<100$): In the short context regime, the system exhibits high volatility.
-This corresponds to a non-equilibrium state in which the information gas has not yet filled the available volume.
-  2. 2.
-    Equilibrium phase ($V_{ctx}>200$): As the volume increases, $\kappa$ converges to a stable constant.
-This stability empirically validates the first law of information dynamics, demonstrating that for a fixed temperature ($T$) and particle number ($N$), the product of pressure and volume remains conserved relative to the system’s efficiency.
+    Data Collapse: Investigate whether scaled fluctuation observables collapse onto a universal curve once the appropriate critical exponents are identified in deeper models.
+VIII.4 Thermodynamic Interpretation of Hallucinations
+We associated hallucinations with thermodynamic fluctuations ($TdS$).
+While thermal noise provides a minimal physical explanation for stochastic errors, hallucinations in large language models likely involve non-equilibrium mechanisms and dataset-driven biases.
+Future studies should incorporate non-equilibrium work relations (e.g., Jarzynski equality) to model the entropy production during the generation of out-of-distribution sequences.
+IX Conclusion
+We have developed an effective thermodynamic description of Transformer attention by formulating its dynamics on an information manifold and constructing a corresponding variational principle.
+Within a Shannon–Boltzmann entropy framework, the Softmax distribution appears as an equilibrium configuration minimizing a free energy functional, suggesting a formal bridge between attention mechanisms and canonical statistical mechanics.
+At the macroscopic level, we introduced a fluctuation-based observable analogous to specific heat and showed empirically that its peak consistently precedes the grokking transition in modular arithmetic tasks.
+Although no asymptotic divergence is observed in shallow, finite systems, the reproducible enhancement of fluctuations indicates a critical-like crossover in the internal energy landscape.
+the results suggest that thermodynamic observables provide a quantitatively testable language for describing representational transitions in neural networks and motivate future investigations into scaling behavior in deeper architectures.